? Fe (ag) reacts with MnO Caq) ion in acidic solution to yield Fe (ag) ions and Mn(a) ions. 5 Fe(aq) + Mno.(aq) + 8H ? 5 Fe(aq) + Mn(aq) + 4H20 A sample of melanterite, FeSO, 7H,O, was analyzed for purity using this reaction by titration of an aqueous solution of the sample. One such sample required 56.15 mL of 0.01301 M permanganate solution to completely titrate all the iron in the sample by the reaction shown above. How much did the sample weigh, if it was in fact pure melanterite? mol melunkvite 1000mL IL B. 0.203 g C. 0.555 g D. 4.851g E. 0.407 g I mo Melo
Answer #1

from equation

5 mol Fe2+ = 1 mol MnO4-

nO of mol of MnO4- required = M*V

                            = 0.01301*56.15

                            = 0.73 mmol

No of mol of Fe2+ reacted = 0.73*5 = 3.65 mmol

   1 mol Fe2+ = 1 mol Feso4.7H2O

No of mol of melanterite reacted (n)= 3.65*10^-3 mol

n = W/M

w = weight of melanterite reacted = ?

M = molarmass of melanterite = 278 g/mol

3.65*10^-3 = w/278

w = weight of melanterite reacted = 1.015 g

answer: A . 1.015 g